.NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE

a)

Brockman Corporation's earnings per share were $3.50 last year, and its growth rate during the prior 5 years was 4.4% per year. If that growth rate were maintained, how many years would it take for Brockman's EPS to triple?

Select the correct answer.

a. 25.51
b. 27.41
c. 19.81
d. 23.61

e. 21.71

b.)

Your friend offers to pay you an annuity of $9,900 at the end of each year for 3 years in return for cash today. You could earn 5.5% on your money in other investments with equal risk. What is the most you should pay for the annuity?

Select the correct answer.

a. $26,709.54
b. $26,713.44
c. $26,701.74
d. $26,705.64

e. $26,717.34

c,)

A new investment opportunity for you is an annuity that pays $2,200 at the beginning of each year for 3 years. You could earn 5.5% on your money in other investments with equal risk. What is the most you should pay for the annuity?

Select the correct answer.

a. $6,279.40
b. $6,261.90
c. $6,296.90
d. $6,314.40
e. $6,244.40

PLEASE ANSWER a.) b.) c.)
Answer #1

(a) Option (a) is correct

Here we will use the following formula:

FV = PV * (1 + r%)n

where, FV = Future value = $3.5 * 3 = $10.5, PV = Present value = $3.5, r = rate of interest = 4.4%, n= time period

now, putting theses values in the above equation, we get,

$10.5 = $3.5 * (1 + 4.4%)n

$10.5 / $3.5 = (1 + 0.044)n

3 = (1.044)n

(1.044)25.51 = (1.044)n

n = 25.51

So, it will take 25.51 years.

(b) Option (a) is correct

Here, the cash inflows will be same every year, so it is an annuity. We will calculate present value of annuity by the following formula:

PVA = P * (1 - (1 + r)-n / r)

where, PVA = Present value of annuity, P is the periodical amount = $9900, r is the rate of interest = 5.5% and n is the time period = 3

Now, putting these values in the above formula, we get,

PVA = $9900 * (1 - (1 + 5.5%)-3 / 5.5%)

PVA = $9900 * (1 - ( 1+ 0.055)-3 / 0.055)

PVA = $9900 * (1 - ( 1.055)-3 / 0.055)

PVA = $9900 * (1 - 0.85161366418) / 0.055)

PVA = $9900 * (0.14838633582 / 0.055)

PVA = $9900 * 2.6979333785

PVA = $26709.54

So, we should pay $26709.54 now for annuity.

(c) Option (b) is correct

Here, the cash inflow will be same every year, so it is an annuity. And since the cash flows will start at the beginning of each year so it will be termed as an annuity due. For calculating the present value of annuity due, we will use the following formula:

PVAD = P * (1 - (1 / (1 + r)n / r) * (1 + r)

where, PVD is the present value of annuity due, P is the periodical amount = $2200, r is the rate of interest = 5.5% and n is the time period = 3

Now, putting these values in the above formula, we get,

PVAD = $2200 * (1 - (1 / (1 + 5.5%)3 / 5.5%) * (1 + 5.5%)

PVAD = $2200 * (1 - (1 / (1 + 0.055)3 / 0.055) * (1 + 0.055)

PVAD = $2200 * (1 - (1 / (1.055)3 / 0.055) * (1.055)

PVAD = $2200 * (1 - (1 / 1.174241375) / 0.055) * (1.055)

PVAD = $2200 * ((1 - 0.8516136641838225) / 0.055) * (1.055)

PVAD = $2200 * (0.1483863358161775 / 0.055) * (1.055)

PVAD = $2200 * 2.697933378475955 * 1.055

PVAD = $6261.90

So, we should pay $6261.90 for this annuity.